# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

### v1

@@ A Balinsky-Novikov structure is given by:
    The Lie superalgebra $D^{12}$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by:
@@
    The Lie superalgebra $D^{13}_{p}$ with $p\not=0$: A Novikov structure is given by:
@@
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
@@
    The Lie superalgebra $D^{14}_{pq}$ with $p\not=0,q\geq 0$: A Novikov structure is given by:
@@
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
@@
    The Lie superalgebra $D^{15}$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
@@
    The Lie superalgebra $D^{16}$: A Novikov structure is given by:
@@
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
@@
    The Lie superalgebra $(A_{3,1}+A)$: A Novikov structure is given by (where $\lambda,\gamma,\mu\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda,\mu\in\mathbb{R})$:
    $\begin{array}[]{lllll}e_{2}\cdot e_{2}=\gamma e_{1},&e_{2}\cdot e_{3}=\lambda e_{1},&e_{3}\cdot e_{2}=(\lambda-1)e_{1},&e_{3}\cdot e_{3}=\mu e_{1}&e_{4}\cdot e_{4}=e_{1}.\end{array}$
@@ A Balinsky-Novikov structure is given by (where $\gamma,\lambda,\mu\in\mathbb{R})$:
    The Lie superalgebra $(D^{3}_{p,-1/2})$ with $p\not=0$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
@@
    The Lie superalgebra $(D^{2}_{-1/2})^{1}$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
@@
    The Lie superalgebra $(D^{2}_{-1/2})^{2}$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
@@
    The Lie superalgebra $(A_{1,1}+3A)^{1}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R}$):
@@ A Balinsky-Novikov structure is given by:
    The Lie superalgebra $(A_{1,1}+3A)^{2}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by:
@@ 6. The list of four-dimensional real Lie superalgebras
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $\omega$ $p(\omega)$
$D^{5}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$D^{6}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{3}]&=&-e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] pq\not=0,\\[2.84526pt] p\geq q\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None
$\begin{array}[]{c}D^{7}_{-1q},\\[2.84526pt] q\leq-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}e_{4}^{*}\wedge e_{1}^{*}+e_{3}^{*}\wedge e_{2}^{*}\end{array}$ ${\bar{1}}$
$\begin{array}[]{c}D^{7}_{pp},\\[5.69054pt] p=-q\\[5.69054pt] q\not=0,-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&-pe_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\end{array}$ $\bar{0}$
$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] p=-q,\\[2.84526pt] q=-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\begin{array}[]{l}\bar{0}\\[2.84526pt] \bar{1}\end{array}$
$\begin{array}[]{c}D^{8}_{p},\\ p\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None
$D^{8}_{-1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}-e_{4}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$\begin{array}[]{c}D^{9}_{pq}\\[2.84526pt] q>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{3}-qe_{4},\\[2.84526pt] [e_{1},e_{2}]&=&e_{2},\\[5.69054pt] [e_{1},e_{4}]&=&qe_{3}+pe_{4}\end{array}$ None
$D^{10}_{q}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&(q+1)e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}(1+q)e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*},\\[5.69054pt] \text{where }q\not=-1\end{array}$ $\bar{1}$
Table 1. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=2|2$
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $p(\omega)$
@@
$\begin{array}[]{c}(D^{7}_{1-p,p})\\[2.84526pt] p\leq\frac{1}{2}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&(1-p)e_{4},\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(D^{8}_{1/2})$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+\frac{1}{2}e_{4},\\[2.84526pt] [e_{4},e_{4}]&=&e_{2}\end{array}$ None --
$(2A_{1,1}+2A)^{1}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
$(2A_{1,1}+2A)^{2}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&e_{1}\\ \end{array}$ None --
$\begin{array}[]{c}(2A_{1,1}+2A)^{3}_{p}\\ p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&p(e_{1}+e_{2})\\ \end{array}$ For p12: NoneFor $\begin{array}[]{l}\text{For $p\not=\frac{1}{2}$: None}\\[2.84526pt] \text{For $p=\frac{1}{2}:\,e_{2}^{*}\wedge e_{3}^{*}-e_{1}^{*}\wedge e_{4}^{*}$}\end{array}$ $\ \bar{1}\end{array}$
$\begin{array}[]{c}(2A_{1,1}+2A)^{4}_{p}\\ p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&p(e_{1}-e_{2})\\ \end{array}$ None --
$(C^{1}_{1}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\\ \end{array}$ $\frac{1}{2}e_{2}^{*}\wedge e_{1}^{*}+\frac{1}{2}e_{3}^{*}\wedge e_{4}^{*}+\frac{1}{4}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(C^{1}_{1/2}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\\ \end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}-\frac{\lambda}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{\mu}{2}e_{4}^{*}\wedge e_{4}^{*},\\ \text{where }\lambda\mu\not=0\end{array}$ $\bar{0}$
$(C^{2}_{-1}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\\ \end{array}$ None --
@@
$(C^{5}_{0}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&-e_{4},&[e_{1},e_{4}]&=&e_{3}\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}&[e_{4},e_{4}]&=&e_{2}\\ \end{array}$ None --
Table 2. Non-trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=2|2$
The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $p(\omega)$
@@
$\begin{array}[]{l}D^{2}_{q}\\[2.84526pt] q\not=-1,0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None --
@@
$\begin{array}[]{l}D^{3}_{pq}\\[2.84526pt] pq\not=0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&pe_{1}-e_{2},&[e_{2},e_{3}]&=&e_{1}+pe_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None --
Table 3. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=3|1$
The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $p(\omega)$
$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|\leq|q|\leq 1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None --
@@
$D^{12}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ None --
$\begin{array}[]{c}D^{13}_{p}\\[2.84526pt] p\not=0\text{ (generic)}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None --
@@
$\begin{array}[]{c}D^{14}_{pq}\\[2.84526pt] p\not=0,q\geq 0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},\\[2.84526pt] [e_{1},e_{3}]&=&qe_{3}-e_{4},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}+e_{3}\end{array}$ None --

### v2

@@ A Balinsky-Novikov structure is given by
    The Lie superalgebra $D^{12}$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure is given by
@@
    The Lie superalgebra $D^{13}_{p}$ with $p\not=0$: A Novikov structure is given by
@@
    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $D^{14}_{pq}$ with $p\not=0,q\geq 0$: A Novikov structure is given by
@@
    A Balinsky-Novikov structure is given by
@@
    The Lie superalgebra $D^{15}$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $D^{16}$: A Novikov structure is given by
@@
    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(A_{3,1}+A)$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{lllll}e_{2}\cdot e_{2}=\gamma e_{1},&e_{2}\cdot e_{3}=\lambda e_{1},&e_{3}\cdot e_{2}=(\lambda-1)e_{1},&e_{3}\cdot e_{3}=\mu e_{1},&e_{4}\cdot e_{4}=e_{1}.\end{array}$
@@ A Balinsky-Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
    The Lie superalgebra $(D^{3}_{p,-1/2})$ with $p\not=0$: A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{2}_{-1/2})^{1}$: A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{2}_{-1/2})^{2}$: A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(A_{1,1}+3A)^{1}$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
@@ A Balinsky-Novikov structure is given by:
    The Lie superalgebra $(A_{1,1}+3A)^{2}$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure is given by
@@
We conclude by saying that the lists of Lie superalgebras classified by Backhouse are all left-symmetric. Moreover, they are all Novikov and Balinsky-Novikov superalgebras, except the two $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$ that are Balinsky-Novikov but not Novikov.
@@ 6. The list of four-dimensional real Lie superalgebras
The tables below differ slightly from those in [BM]. The symplectic forms given in [BM] are given in terms of parameters. For convenience, we specify these parameters here. In addition, we correct the statement regarding the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. The two admit both an odd and an even non-degenerate closed form, contrary to what is stated in [BM].
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $\omega$ $|\omega|$
$D^{5}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$D^{6}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{3},\;[e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{3}]&=&-e_{4},\;[e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] pq\not=0,\\[2.84526pt] p\geq q\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None
$\begin{array}[]{c}D^{7}_{-1q},\\[2.84526pt] q\leq-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}e_{4}^{*}\wedge e_{1}^{*}+e_{3}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$}
$\begin{array}[]{c}D^{7}_{pp},\\[5.69054pt] p=-q\\[5.69054pt] q\not=0,-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&-pe_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\end{array}$ $\bar{0}$
$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] p=-q,\\[2.84526pt] q=-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\begin{array}[]{l}\bar{0}\\[2.84526pt] \bar{1}\end{array}$
$\begin{array}[]{c}D^{8}_{p},\\ p\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None
$D^{8}_{-1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}-e_{4}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$\begin{array}[]{c}D^{9}_{pq}\\[2.84526pt] q>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{3}-qe_{4},&[e_{1},e_{2}]&=&e_{2},\\[5.69054pt] [e_{1},e_{4}]&=&qe_{3}+pe_{4}\end{array}$ None
$D^{10}_{q}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&(q+1)e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4},&[e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}(1+q)e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*},\\[5.69054pt] \text{where }q\not=-1\end{array}$ $\bar{1}$
Table 6. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=2|2$
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $|\omega|$
$(D^{7}_{1/2\;1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{4},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(D^{7}_{1/2\;1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{4},e_{4}]&=&-e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}+\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(D^{7}_{1/2\;1/2})^{3}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{3},e_{3}]&=&e_{2}\end{array}$ None -
$\begin{array}[]{c}(D^{7}_{1-p,p})\\[2.84526pt] p\leq\frac{1}{2}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&(1-p)e_{4},&[e_{3},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(D^{8}_{1/2})$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
$(2A_{1,1}+2A)^{1}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
@@
[…]
@@
[…]
Table 7. Non-trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=2|2$
The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $|\omega|$
@@
$\begin{array}[]{l}D^{2}_{q}\\[2.84526pt] q\not=-1,0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None -
@@
$\begin{array}[]{l}D^{3}_{pq}\\[2.84526pt] pq\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{1}-e_{2},&[e_{2},e_{3}]&=&e_{1}+pe_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None -
Table 8. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=3|1$
The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $|\omega|$
$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|\leq|q|\leq 1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None -
@@
$D^{12}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ None -
$\begin{array}[]{c}D^{13}_{p}\\[2.84526pt] p\not=0\text{ (generic)}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None -
@@
$\begin{array}[]{c}D^{14}_{pq}\\[2.84526pt] p\not=0,q\geq 0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},\\[2.84526pt] [e_{1},e_{3}]&=&qe_{3}-e_{4},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}+e_{3}\end{array}$ None -